OpenStudy (anonymous):

solve for t. e^1=t/t-5

OpenStudy (anonymous):

if u can just replace e^1 with 2.71 (which is e), then it'll be 2.71=t/(t-5). 2.71t - 13.591 = t then t= 7.94

OpenStudy (anonymous):

or does your teacher want you to keep the e?

OpenStudy (anonymous):

it is just asking to solve for t. the problem is longer than that, but i managed to simplify it, hopefully right. the problem was: ln(t) - ln(t-5) = 1

OpenStudy (anonymous):

i believe you're right

OpenStudy (anonymous):

okay. lets cross our fingers. I have one more try left.

OpenStudy (anonymous):

no. the answer was t= 5 x e/(e-1)

OpenStudy (anonymous):

thanks for the help tho!!

OpenStudy (anonymous):

srry i dont know how to solve it :(

OpenStudy (anonymous):

its okay!!!

OpenStudy (anonymous):

ln (t) - ln(t-5) = 1 ln[t/(t-5)] =1 t/(t-5) = e t = e t - 5e t(e-1) = 5e t = 5e/(e-1)